I need help with this problem please.
f(x)=5x+3; g(x)=x² and solve for (g0g)(x)
Thank you so much

Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

f(x)=5x+3; g(x)=x² and solve for (g0g)(x)
------------
g(x) = x^2
Find g(x^2)
Sub x^2 for x in g(x)

Answer by fractalier(6550)

(g o g)(x) means g(g(x)) = g(x^2) = x^4